There are 200 people in the population. Ten of them are under the age of 16 and have no job. Five of them are under the age of 16 and have a job. Thirty people have retired, and half of these are no longer capable of working. There are 25 stay at home spouses. There are 20 college students with no job and half of them are looking for a job. Ten people have no job but are looking, and 10 people have no job but are not looking.

The rest of the people have jobs.

There are ? people in the adult (working age) population and ? people in the labor force. Enter whole numbers.

The unemployment rate is ? % and the labor force participation rate is ? %. Round to the nearest whole number.
Answer #1

Number of people in adult working age population are the retired, the stay at home spouses, the college students, people with no jobs and looking for one and people with no jobs and not looking for one. = 30 + 25 + 20 + 10 + 10 = 95

Number of people in the labor force are people in adult working population minus the people not looking for a job = 95 - 15 - 25 = 55

Unemployment rate is = number of unemployed members of the labor force/ number of people in the labor force = 10 + 10 / 55 = 0.36